Pawlak and topological rough sets in terms of multifunctions

P. Maritz, Stellenbosch, South Africa

AMS Subject classification: 54C60, 54A99, 04A99

Keywords: approximation space, multifunction, rough sets

Abstract: The purpose of this paper is to discuss approximation spaces in terms of surjective multifunctions. These multifunctions will be mainly semi-single-valued. The topologies associated with the upper and lower approximation operations will be studied.